OpenStudy (kkbrookly):

Evaluate cos2a if sin a=3/5, pi/2

jimthompson5910 (jim_thompson5910):

http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf see page 2 (double angle formulas) \[\Large \cos(2a) = 1 - 2\sin^2(a)\] \[\Large \cos(2a) = 1 - 2*\left(\sin(a)\right)^2\] \[\Large \cos(2a) = 1 - 2*\left(\frac{3}{5}\right)^2\] I'll let you finish up

jimthompson5910 (jim_thompson5910):

Special note: if pi/2<a<pi then pi<2a<2pi notice how I doubled everything

OpenStudy (kkbrookly):

cos(2a)=7/25, right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (kkbrookly):

Do I do anything to the cos(2a)? Or is that the answer?

OpenStudy (kkbrookly):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

nope, cos(2a) is left alone and yes that is the answer

OpenStudy (kkbrookly):

Thank you!! Can you help me with another one?

jimthompson5910 (jim_thompson5910):

sure